# Contrasting Features of Parton Energy Loss in Heavy-ion Collisions at RHIC and the LHC

### v3

@@ Abstract
Energetic quarks and gluons lose energy as they traverse the hot and dense medium created in high-energy heavy-ion collisions at the BNL Relativistic Heavy Ion Collider (RHIC) and the CERN Large Hadron Collider (LHC). The nuclear modification factor ($R_{AA}$) of leading particles quantifies parton energy loss in such collisions, with the particle spectrum in $p+p$ collisions as a reference. Previous $R_{AA}$ measurements at RHIC energies have revealed an approximately constant trend at high transverse momenta ($p_{T}$), implying a scenario where parton energy loss, $\Delta p_{T}$, scales proportionally with $p_{T}$, a feature naively expected from energy loss dynamics in elastic collisions. In this study, we investigate the LHC $R_{AA}$ measurements which exhibit a pronounced $p_{T}$ dependence of $R_{AA}$ for various particle species, and our analysis attributes this behavior to $\Delta p_{T}$ being approximately proportional to $\sqrt{p_{T}}$. These distinct features are consistent with model calculations of dominant radiative energy loss dynamics at the LHC, in contrast to the dominance of collisional energy loss at RHIC. Additionally, the linear increase of fractional energy loss with medium density at different $p_{T}$ magnitudes affirms the previous empirical observation that the magnitude of the energy loss depends mostly on the initial entropy density, with no significant path-length dependence. Implications on the dynamical scenarios of parton energy loss and future experimental investigations will also be discussed.
@@
Color opacity stands as a fundamental trait of the hot and dense medium created in heavy-ion collisions at the BNL Relativistic Heavy Ion Collider (RHIC) and the CERN Large Hadron Collider (LHC). As energetic quarks and gluons traverse the medium, they shed energy through elastic scattering [1, 2, 3, 4] and radiation of soft gluons [5, 6, 7].
@@
In the scenario of an infinitely-high-momentum parton or an infinitely massive scattering center, energy loss would predominantly occur through radiative processes. Conversely, in the opposite scenario, collisional energy loss would become the dominant factor. Prior empirical examinations of final-state leading particle spectra and the pertinent nuclear effects, using RHIC data, have revealed the proportionality between parton energy loss ($\Delta p_{T}$) and the magnitude of transverse momentum ($p_{T}$[8, 9, 10]. Although most theoretical energy loss calculations at RHIC are framed through the lens of radiative energy loss [11, 12], a proportional relationship between parton energy loss and transverse momentum tends to correspond better with a “classical mechanics” understanding of elastic scattering and collisional energy loss.
@@
Given that collision center-of-mass energies ($\sqrt{s_{NN}}$) at the LHC significantly surpass those at RHIC by over an order of magnitude, the associated $p_{T}$ range of generated particles now spans into a realm where radiative energy loss dynamics are expected to assume a more prominent role. Parton energy loss in the LHC center-of-mass energy regime has been previously studied through analysis of jet suppression $S_{\text{loss}}$ results with ATLAS [13], but performing a similar study using leading particles instead allows for a cleaner look into the nature of the energy loss without some of the compounding effects within jets. Hence, the analysis of LHC data using the same framework as in Ref. [8] is warranted to investigate the potential transition in the dynamics of energy loss from RHIC to the LHC through the study of leading particle data instead of jet data.
@@
Both radiative and collisional energy losses are intricately linked to the path length ($L$) and the entropy density of the medium. We approximate the medium entropy density as $\frac{1}{S}\frac{dN}{dy}$, where $\frac{dN}{dy}$ represents the experimentally measured particle density per unit rapidity, and $S$ corresponds to the transverse overlap area of the colliding system, which can be determined using Monte Carlo Glauber calculations [14, 15, 16, 17]. A previous study of RHIC data has unraveled the lack of or minimal dependence of $\Delta p_{T}$ on $L$, implying that parton energy loss is predominantly determined by the initial medium density [8]. This feature could arise from the scenario of rapid expansion of the collision system, resulting in a swift decrease in medium entropy density over time. It is of great interest to investigate whether the LHC data corroborate the same characteristics.
@@
To recap, the analyzed LHC data here suggest that to explain the $R_{AA}$ measurements for light- and heavy-quark hadrons as a $p_{T}$ shift in the spectrum from $p$+$p$ collisions, we require the corresponding $\Delta p_{T}$ to scale with $\sqrt{p_{T}}$. This $p_{T}$ dependence contrasts with the previously observed proportionality with $p_{T}$ in RHIC data. Our analysis results with more recent data are in line with a previous study of LHC $R_{AA}$ data that determined the $\alpha$ value to be 0.55 [31]. Our $p_{T}$ dependence of the parton energy loss at LHC supports theoretical predictions involving energy loss dynamics from medium-induced gluon radiation [32]. The distinct change from $\alpha=1$ at RHIC to $\alpha=0.5$ at LHC suggests a transition in the relative importance of collisional energy loss dynamics to radiative energy loss dynamics.
This transition in the parton energy loss dynamics within the medium might also find an explanation in the effects of the jet-medium running coupling constant. Theoretical treatment of running coupling effects in both the radiative and elastic contributions to energy loss were previously shown to have robust agreement with LHC and RHIC data for light- and heavy-flavor $R_{AA}$ measurements [33]. In this investigation, the emphasis was placed on the contributions from the radiated gluon vertex to the DGLV integral, particularly noting their significant role in enhancing the color transparency of the medium for higher-energy jets [33]. Should this apply to leading particles in addition to jets, our conclusion that radiative energy loss has an increased importance at the LHC relative to RHIC would be corroborated.
@@
The dead-cone effect [34] predicts that gluon radiation is more strongly suppressed for bottom quarks than charm quarks, as the former bears a larger mass-to-energy ratio, leading to a wider dead cone. Recent measurements of heavy-quark meson production in $p$+$p$ collisions by the ALICE experiment [35] reveal heavy-quark fragmentation in the vacuum and provide a direct observation of the dead-cone effect. However, the LHC data of the $R_{AA}$ trends for muons from charm and bottom decays do not exhibit the anticipated reduced radiative energy loss for bottom quarks. The decay muon measurements can be influenced by various factors, including substantial momentum smearing resulting from decay kinematics and the existence of non-prompt $c\rightarrow\mu$ decays that originate from $b$ quarks. Another factor to consider is that the dead-cone effect may become less pronounced for very-high-energy quarks represented in the muon measurements. More precise data are needed to elucidate the nature of heavy-quark dynamics in the medium.
@@
In their similar study of $S_{\text{loss}}$, ATLAS additionally investigated the apparent differences in quark and gluon jet suppression through comparing inclusive and photon-tagged jet spectra [13]. While we do not explore this relationship in our study due to the challenging task of identifying whether our leading particle originates from a quark or gluon, it is crucial to underscore these prior findings and acknowledge their significance in the context of parton energy loss.
@@
We do, however, also investigate the relationship between energy loss and path length at LHC energies. Previous examinations of RHIC data have revealed that the deduced fractional energy loss, $\Delta p_{T}/p_{T}$, is a linear function of medium initial entropy density (quantified by $\frac{1}{S}\frac{dN}{dy}$) across different centrality intervals, despite significant variations in the path length for traversing partons [8].
@@
This suggests a weak dependence of energy loss on path length and instead puts a stronger emphasis on initial collision density. We apply the same analysis to LHC data, and discover similar outcomes, as shown in Fig. 6. Now that fractional energy loss varies with $p_{T}$ according to the LHC $R_{AA}$ data, we plot $\Delta p_{T}/p_{T}$ as a function of $\frac{1}{S}\frac{dN}{dy}$ at different $p_{T}$ values for charged hadrons in Pb+Pb collisions at 2.76 TeV and for charged pions in Pb+Pb collisions at 5.02 TeV. In each case for each $p_{T}$ regime, a clear linear trend emerges, and the linearity is especially strong for higher $p_{T}$ scales, where parton fragmentation dominates particle production [37].
@@
We also find that the fractional energy loss decreases with increased transverse momentum at LHC energies, agreeing with the trend observed in previous ATLAS $S_{\text{loss}}$ measurements for inclusive jets [13], as well as those found in Ref. [37]. The linear trends between $\Delta p_{T}/p_{T}$ and $\frac{1}{S}\frac{dN}{dy}$ support the earlier observation of a weak path-length dependence of parton energy loss, even though the medium densities at RHIC and the LHC are very different. As discussed for RHIC data [8], the weak path-length dependence of energy loss could arise from the rapid expansion of the medium, where the majority of energy loss occurs before the parton is able to traverse a full path length. Thereby, medium density becomes the dominant factor that determines the energy loss during the rapid expansion. We argue that in such a rapidly expansive medium, the static path length from the initial geometry of colliding nuclei fails to be a dominating factor for the parton energy loss in the medium. Our parton energy loss scenario would predict that the elliptic flow of jets or leading high $p_{T}$ particles in A+A collisions of medium centrality, if measured with minimum non-flow correlation effect, will be close to zero, much smaller than what may be calculated from the static path length dependence.
@@
With this observation and prediction, it becomes more difficult to explain the large $v_{2}$ measurements observed for jets and leading particles at high $p_{T}$ [38]. Previous calculations of nuclear modification factor and anisotropic flow values for jets at RHIC and the LHC, however, do not prevent this possibility. In Ref. [39], a wide variety of energy loss models are shown to be able to describe available RHIC and LHC data within experimental uncertainty for both strong and weak path-length-dependent models. While this study indicates that the absence of path-length dependence is disfavored in the explored models inspired by quantum chromodynamics, it does not constrain the path-length-dependent parameter to be larger than zero [39]. The exact degree of path-length dependence of parton energy loss remains an outstanding physics issue, which should be addressed with more precise data. The nonflow correlations in elliptic flow measurements for jets and high-$p_{T}$ particles present an important factor to be accounted for as well.
@@
In summary, we present a parton energy loss study showing a significant distinction between RHIC and LHC data when empirically interpreting $R_{AA}(p_{T})$ as a momentum loss in $A$+$A$ collisions relative to the $p$+$p$ reference. While the RHIC data favor a direct proportionality between the $p_{T}$ shift and $p_{T}$ itself, the LHC data suggest a proportionality with $\sqrt{p_{T}}$. This difference in the $p_{T}$ dependence signifies the heightened importance of radiative energy loss compared with collisional energy loss within the same transverse momentum range in colliding systems at higher $\sqrt{s_{NN}}$. Additionally, we find that the magnitude of the parton energy loss at LHC is largely determined by the initial medium entropy density, consistent with previous results at RHIC, indicating a limited path-length dependence of parton energy loss, and placing greater emphasis on the initial medium density for a rapid explosive medium. The distinct parton energy loss dynamics at RHIC and the LHC can be further investigated with high-statistics heavy-quark-tagged jets from the sPHENIX experiment at RHIC, as well as the LHC experiments in future runs. Equally interesting is the differentiation of energy loss contributions from quarks and gluons to the measurements studied here.

### v4

@@ Abstract
Energetic quarks and gluons lose energy as they traverse the hot and dense medium created in high-energy heavy-ion collisions at the BNL Relativistic Heavy Ion Collider (RHIC) and the CERN Large Hadron Collider (LHC). The nuclear modification factor ($R_{AA}$) of leading particles quantifies parton energy loss in such collisions, with the particle spectrum in $p+p$ collisions as a reference. Previous $R_{AA}$ measurements at RHIC energies have revealed an approximately constant trend at high transverse momenta ($p_{T}$), implying a scenario where parton energy loss, $\Delta p_{T}$, scales proportionally with $p_{T}$, a feature naively expected from energy loss dynamics in elastic collisions. In this study, we investigate the LHC $R_{AA}$ measurements which exhibit a pronounced $p_{T}$ dependence of $R_{AA}$ for various particle species, and our analysis attributes this behavior to $\Delta p_{T}$ being approximately proportional to $\sqrt{p_{T}}$. These distinct features are consistent with model calculations of dominant radiative energy loss dynamics at the LHC, in contrast to the dominance of collisional energy loss at RHIC. Additionally, the linear increase of the fractional energy loss with medium density at different $p_{T}$ magnitudes affirms our previous empirical observation that the magnitude of the energy loss depends more strongly on the initial entropy density rather than the parton’s path length through the medium. Implications on the dynamical scenarios of parton energy loss and future experimental investigations will also be discussed.
@@
Color opacity stands as a fundamental trait of the hot and dense medium created in heavy-ion collisions at the BNL Relativistic Heavy Ion Collider (RHIC) and the CERN Large Hadron Collider (LHC). As energetic quarks and gluons traverse the medium, they shed energy through elastic scattering [1, 2, 3, 4] and radiation of soft gluons [5, 6, 7]. In the scenario of an infinitely-high-momentum parton or an infinitely massive scattering center, energy loss would predominantly occur through radiative processes. Conversely, in the opposite scenario, collisional energy loss would become the dominant factor. Prior empirical examinations of final-state leading particle spectra and the pertinent nuclear effects, using RHIC data, have revealed the proportionality between parton energy loss ($\Delta p_{T}$) and the magnitude of transverse momentum ($p_{T}$[8, 9, 10]. Although most theoretical energy loss calculations at RHIC are framed through the lens of radiative energy loss [11, 12], a proportional relationship between parton energy loss and transverse momentum tends to correspond better with a “classical mechanics” understanding of elastic scattering and collisional energy loss, where higher-momentum bodies tend to lose proportionally more energy in collisions. Previous theoretical predictions additionally seem to show a linear dependence of the energy loss on the energy magnitude for heavy quarks, which would align with a similar dependence on $p_{T}$ [13].
@@
Given that collision center-of-mass energies ($\sqrt{s_{NN}}$) at the LHC significantly surpass those at RHIC by over an order of magnitude, the associated $p_{T}$ range of generated particles now spans into a realm where radiative energy loss dynamics are expected to assume a more prominent role. Parton energy loss in the LHC center-of-mass energy regime has been previously studied through analysis of jet suppression $S_{\text{loss}}$ results with ATLAS [14], but performing a similar study using leading particles instead allows for a cleaner look into the magnitude and the $p_{T}$ dependence of the energy loss without some compounding effects within jets. If energy loss and fragmentation were to occur sequentially—where fragmentation takes place outside the medium after the leading parton loses energy within the medium—leading particles would provide a clearer means of investigation than jets. Recent theoretical simulations of jet-induced medium excitation examine the diffusion wake in high-$p_{T}$ jet-photon pairs and predict an identical width, both with and without accounting for the medium, suggesting the possibility of separating energy loss and fragmentation [15]. Hence, the analysis of LHC data using the same framework as in Ref. [8] is warranted to investigate the potential transition in the dynamics of energy loss from RHIC to the LHC through the study of leading particle data instead of jet data.
@@
Both radiative and collisional energy losses are intricately linked to the path length ($L$) and the entropy density of the medium. We approximate the medium entropy density as $\frac{1}{S}\frac{dN}{dy}$, where $\frac{dN}{dy}$ represents the experimentally measured particle density per unit rapidity, and $S$ corresponds to the transverse overlap area of the colliding system, which can be determined using Monte Carlo Glauber calculations [16, 17, 18, 19]. A previous study of RHIC data has unraveled a minimal dependence of $\Delta p_{T}$ on $L$, implying that parton energy loss is predominantly determined by the initial medium density [8]. This feature could arise from the scenario of rapid expansion of the collision system, resulting in a swift decrease in medium entropy density over time. It is of great interest to investigate whether the LHC data corroborate the same characteristics.
@@
To recap, the analyzed LHC data here suggest that to explain the $R_{AA}$ measurements for light- and heavy-quark hadrons as a $p_{T}$ shift in the spectrum from $p$+$p$ collisions, we require the corresponding $\Delta p_{T}$ to scale with $\sqrt{p_{T}}$. This $p_{T}$ dependence contrasts with the previously observed proportionality with $p_{T}$ in RHIC data. Our analysis results with more recent data are in line with a previous study of LHC $R_{AA}$ data that determined the $\alpha$ value to be 0.55 [33]. Our $p_{T}$ dependence of the parton energy loss at LHC supports theoretical predictions involving energy loss dynamics from medium-induced gluon radiation [34]. Thus, the distinct change from $\alpha=1$ at RHIC to $\alpha=0.5$ at LHC suggests a transition in the relative importance of collisional energy loss dynamics to radiative energy loss dynamics.
[…]
@@
The dead-cone effect [36] predicts that gluon radiation is more strongly suppressed for bottom quarks than charm quarks, as the former bears a larger mass-to-energy ratio, leading to a wider dead cone. Recent measurements of heavy-quark meson production in $p$+$p$ collisions by the ALICE experiment [37] reveal heavy-quark fragmentation in the vacuum and provide a direct observation of the dead-cone effect. However, the LHC data of the $R_{AA}$ trends for muons from charm and bottom decays do not exhibit the anticipated reduced radiative energy loss for bottom quarks. The decay muon measurements can be influenced by various factors, including substantial momentum smearing resulting from decay kinematics, reduced sensitivity to low-momentum heavy-quark mesons, and the existence of non-prompt $c\rightarrow\mu$ decays that originate from $b$ quarks. Recent CMS [38] and ALICE [39] results indicate a significant reduction in the suppression of non-prompt charm hadrons over prompt charm hadrons, which could, in particular, contribute to the differences we see here via the latter factor. Another factor to consider is that the dead-cone effect may become less pronounced for very-high-energy quarks represented in the muon measurements. More precise data are needed to elucidate the nature of heavy-quark dynamics in the medium.
@@
In their similar study of $S_{\text{loss}}$, ATLAS additionally investigated the apparent differences in quark and gluon jet suppression through comparing inclusive and photon-tagged jet spectra [14]. These differences were also explored in recent studies of the $p_{T}$ dispersion of inclusive jets and production cross sections of $J/\Psi$ mesons [40, 41]. While we do not explore this relationship in our study due to the challenging task of identifying whether our leading particle originates from a quark or gluon, it is crucial to underscore these prior findings and acknowledge their significance in the context of parton energy loss.
@@
We do, however, also investigate the relationship between energy loss and path length at LHC energies. Previous examinations of RHIC data have revealed that the deduced fractional energy loss, $\Delta p_{T}/p_{T}$, is a linear function of medium initial entropy density (quantified by $\frac{1}{S}\frac{dN}{dy}$) across different centrality intervals, despite significant variations in the path length for traversing partons [8]. This suggests that the initial collision density plays a more significant role in determining parton energy loss than the path length traversed by those partons. We apply the same analysis as adopted in Ref. [8] to the LHC data and discover similar outcomes, as shown in Fig. 6. Now that fractional energy loss varies with $p_{T}$ according to the LHC $R_{AA}$ data, we plot $\Delta p_{T}/p_{T}$ as a function of $\frac{1}{S}\frac{dN}{dy}$ at different $p_{T}$ values for charged hadrons in Pb+Pb collisions at 2.76 TeV and for charged pions in Pb+Pb collisions at 5.02 TeV. In each case for each $p_{T}$ regime, a clear linear trend emerges, and the linearity is especially strong for higher $p_{T}$ scales, where parton fragmentation dominates particle production [43].
@@
We also find that the fractional energy loss decreases with increased transverse momentum at LHC energies, agreeing with the trend observed in previous ATLAS $S_{\text{loss}}$ measurements for inclusive jets [14], as well as those found in Ref. [43]. The linear trends between $\Delta p_{T}/p_{T}$ and $\frac{1}{S}\frac{dN}{dy}$ support the previous observation that the initial collision density has a stronger impact than path length, despite the significantly different initial densities at RHIC and the LHC. As discussed regarding RHIC data [8], a subdominant path-length dependence of energy loss might result from the rapid expansion of the medium, where most energy loss takes place before the parton traverses the full path length. Thereby, medium density becomes the dominant factor that determines the energy loss during the rapid expansion. Recent theoretical studies also suggest the formation time of partons could play a strong role in determining the medium induced energy loss [44]. We argue that in such a rapidly expansive medium, the static path length from the initial geometry of colliding nuclei fails to be the dominating factor for the parton energy loss in the medium.
@@
Our parton energy loss scenario predicts that the elliptic flow ($v_{2}$) of jets or high-$p_{T}$ leading particles in A+A collisions of intermediate centrality, if measured with minimal nonflow effects, will be close to zero and significantly smaller than values calculated assuming a static path-length dependence. Therefore, it is challenging to explain the large $v_{2}$ values observed for jets and leading particles at high $p_{T}$ [45]. Previous calculations of nuclear modification factor and anisotropic flow values for jets at RHIC and the LHC, however, do not prevent this possibility. In Ref. [46], a wide variety of energy loss models are shown to be able to describe available RHIC and LHC data within experimental uncertainty for both strong and weak path-length-dependent models. While this study indicates that the absence of path-length dependence is disfavored in the explored models inspired by quantum chromodynamics, it does not constrain the path-length-dependent parameter to be larger than zero [46]. The exact degree of path-length dependence of parton energy loss remains an outstanding physics issue, which should be addressed with more precise data. It is puzzling why the observed $v_{2}$ for jets or leading particles shows a weak $p_{T}$ dependence, while the magnitude of parton energy loss exhibits a strong $p_{T}$ dependence. There are some experimental details worthy of more investigation such as the subtraction scheme for the $v_{2}$-dependent background. The nonflow correlations in elliptic flow measurements for jets and high-$p_{T}$ particles present an important factor to be accounted for as well. These experimental concerns should be adequately addressed before the path length dependence of the parton energy loss may be definitively determined.
@@
In summary, we present a parton energy loss study showing a significant distinction between RHIC and LHC data when empirically interpreting $R_{AA}(p_{T})$ as a momentum loss in $A$+$A$ collisions relative to the $p$+$p$ reference. While the RHIC data favor a direct proportionality between the $p_{T}$ shift and $p_{T}$ itself, the LHC data suggest a proportionality with $\sqrt{p_{T}}$. This difference in the $p_{T}$ dependence signifies the heightened importance of radiative energy loss compared with collisional energy loss within the same transverse momentum range in colliding systems at higher $\sqrt{s_{NN}}$. Additionally, we find that the magnitude of the parton energy loss at LHC is largely determined by the initial medium entropy density, consistent with previous results at RHIC. This indicates that the path-length dependence of parton energy loss is less dominant, placing a greater emphasis on the initial medium density for a rapidly expanding medium. The distinct parton energy loss dynamics at RHIC and the LHC can be further investigated with high-statistics heavy-quark-tagged jets from the sPHENIX experiment at RHIC, as well as the LHC experiments in future runs. The possible sequential treatment of fragmentation and parton energy loss is of great interest for these future experiments. Equally interesting is the differentiation of energy loss contributions from quarks and gluons to the measurements studied here.